OpenStudy (anonymous):

Evaluate f(x)=x^2+1 for f(10)

OpenStudy (midhun.madhu1987):

Replace x by 10 and find the value...

OpenStudy (anonymous):

What like 10^2+1=

OpenStudy (midhun.madhu1987):

yes...

OpenStudy (anonymous):

So 21 right

OpenStudy (zale101):

no, this is what @midhun.madhu1987 meant f(x)=x^2+1 what is the equation doing to x? It's squaring x and then adding 1, so do that with f(10) not f(x). What is the function doing to 10? It's squaring it and adding 1. \[f(10)=(10)^2+1\] now solve

OpenStudy (anonymous):

So is it 21

OpenStudy (midhun.madhu1987):